Condensed financial data of Fairfield Company for 2010 and 2009 are presented below:

Comparative Balance Sheet

as at December 31, 2010 and 2009

2010

2009

Cash

$2,150

$1,110

Receivables

1,750

1,300

Inventory

1,600

1,900

Plant Assets

1,910

1,700

Accumulated Depreciation

(1,200)

(1,170)

Long-term investments (Held-to-Maturity)

1,300

1,470

$7,510

$6,310

Accounts Payable

$1,250

$800

Income Tax Payable

Accrued Liabilities

      90   

200

   10  

250

Bonds Payable

1,400

1,650

Capital Stock

1,910

1,700

Retained Earnings

2,660

1,900

$7,510

$6,310

Income Statement

For the year ended December 31, 2010

Sales

$7,000

Cost of goods sold

4,700

Gross margin

2,300

Depreciation

Selling and administrative expense

30

900

Income from operations

1,370

Loss on sale of investments

     70

Income before tax

1,300

Income tax expense

   390

Net Income

   910

Additional information: No plant assets were sold in 2010. Cash dividends were $150.

Required:

Prepare a statement of cash flows for 2010 using the indirect method.

Using the direct method, prepare the Operating Activities section of the statement of cash flows for 2010.                                                                                
Answer #1

Solution:

Cash flow statement using indirect method:

Workings:

Payments to accounts payable = Beg. accounts payble + purchase - Ending accounts payable

= $800 +(4700+1600-1900) - 1250 = $3,950

Amount received from accounts receivable = beg. balance + sales - ending bal.

Payment of income tax = Beg. bal of income tax payable + income tax exp. - Ending bal. of income tax payable

= 10 + 390 - 90 =$310